OpenStudy (anonymous):

find height of a tree that makes a 40'shadow when5'3" in pole is used to make a shadow of 3'

OpenStudy (anonymous):

we have..... 3 40 ---=---- 5'3" h h=40(5+3/12)/3 = 40*21/12 =70'

OpenStudy (anonymous):